Jefferson Lab HALL A Logbook for September 08 - All EXCEPT Automatic Entries

  • Main INDEX

    Tuesday

  • 243468 09/30/08 23:00 adev BigBite WireChamber Measured Voltage change!
  • 243464 09/30/08 22:49 adev BigBite WireChamber Measured Voltage change!
  • 243460 09/30/08 22:29 adev BigBite WireChamber Threshold change!
  • 243440 09/30/08 15:02 adev BigBite WireChamber Threshold change!
  • 243437 09/30/08 14:21 Kalyan HV set for run 1632
  • Monday

  • 243426 09/29/08 14:30 Moffit F1 tdc change..
  • 243421 09/29/08 10:31 Yi Qiang Water NMR Calibration
  • Sunday

  • 243419 09/28/08 17:13 adev BigBite WireChamber HVGUI Start!
  • Friday

  • 243417 09/26/08 16:18 Jin Huang Target: Verticle Coil Polarity Inversed
  • 243416 09/26/08 11:27 adev BigBite WireChamber Measured Voltage change!
  • 243415 09/26/08 09:28 R. Michaels Right HRS scalers dropped
  • 243414 09/26/08 09:08 V. Sulkosky Rebooted IOCHA5 for BCM
  • Thursday

  • 243413 09/25/08 23:43 Yi Qiang First Water Signal in Hall !
  • 243410 09/25/08 15:37 adev BigBite WireChamber HVGUI Start!
  • 243409 09/25/08 15:16 Kalyan BigBite DAQ is up and running again
  • 243407 09/25/08 15:08 kalyan Compressed air lines connected for BigBite PMTs
  • Wednesday

  • 243405 09/24/08 12:33 Kalyan CPU borads for upgrade to Vxworks 5.5
  • 243404 09/24/08 09:47 K DAQ down
  • Tuesday

  • 243401 09/23/08 16:57 adev BigBite WireChamber Threshold change!
  • 243394 09/23/08 16:19 V. Sulkosky Ped Suppression Removed for Helicity and Target Spin
  • 243393 09/23/08 16:15 adev BigBite WireChamber Threshold change!
  • 243392 09/23/08 15:54 adev BigBite WireChamber Threshold change!
  • 243391 09/23/08 15:13 adev BigBite WireChamber HVGUI Start!
  • Monday

  • 243387 09/22/08 18:13 adev BigBite WireChamber HVGUI Start!
  • 243386 09/22/08 17:55 Kalyan Cosmic trigger for BB lead-glass blocks
  • Sunday

  • 243380 09/21/08 16:17 Kalyan bbts1 crate in the readout as ROC12
  • Saturday

  • 243377 09/20/08 21:19 Jack VDC gas flowing
  • Friday

  • 243366 09/19/08 17:25 Jin Huang RICH lib added to replay softwares
  • 243355 09/19/08 10:54 R. Michaels R-HRS DAQ ready, runs 2.6
  • Thursday

  • 243348 09/18/08 16:19 xqian Noise Cross-talk on BigBite Wire Chamber
  • 243341 09/18/08 15:38 Jack Left Arm Detector Stack is out
  • 243334 09/18/08 14:37 Moffit Left Arm TIR/TS cable disconnected
  • 243326 09/18/08 11:45 R. Michaels Modified ts_scaler.crl (LHRS)
  • 243324 09/18/08 11:19 Jack Dry air to the hall
  • Wednesday

  • 243315 09/17/08 21:20 V. Sulkosky Target Spin Signal Added to TIR
  • 243314 09/17/08 21:14 V. Sulkosky Helicity and Target Spin Signal in ADC
  • 243295 09/17/08 14:13 R. Michaels scaler swap on L-HRS
  • Tuesday

  • 243285 09/16/08 23:41 adev BigBite WireChamber Measured Voltage change!
  • 243284 09/16/08 23:31 adev BigBite WireChamber Measured Voltage change!
  • 243282 09/16/08 22:28 V. Sulkosky Re: Scaler gates and target signals in L-HRS
  • 0 09/16/08 18:00 no_name no_key
  • 243263 09/16/08 14:08 Jin Huang Target: Spin Signal to IOC changed to TTL level
  • 243258 09/16/08 12:15 adev BigBite WireChamber Measured Voltage change!
  • 243255 09/16/08 12:03 adev BigBite WireChamber Measured Voltage change!
  • 243245 09/16/08 11:11 adev BigBite WireChamber Measured Voltage change!
  • Monday

  • 243242 09/15/08 23:10 adev BigBite WireChamber Measured Voltage change!
  • 243232 09/15/08 21:47 adev BigBite WireChamber Measured Voltage change!
  • 243211 09/15/08 18:46 Kalyan Scaler gates and target signals in L-HRS
  • 243210 09/15/08 18:45 Kalyan Scaler gates and target signals in L-HRS
  • 243203 09/15/08 17:11 Jin Huang hareboot12 set up for In Hall Target electronics
  • 243200 09/15/08 15:56 brads Outside web access blocked for computers on Hall A subnets
  • 243199 09/15/08 15:44 Jin Huang Proxy Setting to Access hareboot*
  • 243191 09/15/08 10:16 adev BigBite WireChamber HVGUI Start!
  • 243190 09/15/08 09:31 ycwang A1 information of cosmic test
  • 243188 09/15/08 09:15 adev BigBite WireChamber Measured Voltage change!
  • 243186 09/15/08 09:05 adev BigBite WireChamber Measured Voltage change!
  • Sunday

  • 243185 09/14/08 23:53 adev BigBite WireChamber Measured Voltage change!
  • 243184 09/14/08 23:43 adev BigBite WireChamber Threshold change!
  • 243183 09/14/08 23:32 adev BigBite WireChamber Measured Voltage change!
  • 243182 09/14/08 18:57 V. Sulkosky Turned on LHRS VDC Discriminator Thresholds
  • 243172 09/14/08 16:58 adev BigBite WireChamber Threshold change!
  • 243159 09/14/08 13:51 adev BigBite WireChamber Threshold change!
  • 243158 09/14/08 13:41 adev BigBite WireChamber Threshold change!
  • Saturday

    
    

    Friday

  • 243145 09/12/08 22:38 adev BigBite WireChamber Threshold change!
  • 243142 09/12/08 19:05 V. Sulkosky Re: S2m ADC Signals from Paddle 4 Resolved
  • 243141 09/12/08 18:45 V. Sulkosky ADC in ROC 4 Slot 20 Removed
  • Thursday

  • 243073 09/11/08 16:06 adev BigBite WireChamber Threshold change!
  • 243072 09/11/08 15:51 Jin Huang Target: Spin Signal Recovered
  • 243069 09/11/08 14:54 Jin Huang Target: Spin Signal to Bigbite to be down for 1 hour
  • 243065 09/11/08 14:22 R. Michaels adaql2 rebuilt, CODA running
  • 243063 09/11/08 14:12 adev BigBite WireChamber Threshold change!
  • Wednesday

  • 243062 09/10/08 18:22 adev BigBite WireChamber Measured Voltage change!
  • 243061 09/10/08 18:11 adev BigBite WireChamber Measured Voltage change!
  • 243060 09/10/08 18:11 V. Sulkosky A1 Aerogel Discriminator Threshold Changed (30 mV)
  • 243059 09/10/08 18:05 V. Sulkosky Re: L-HRS DAQ problems fixed
  • 243058 09/10/08 17:42 V. Sulkosky S2m ADC Signals from Paddle 4 Resolved(?)
  • 243057 09/10/08 14:46 atrig Trigger-Download
  • 243054 09/10/08 01:03 adev BigBite WireChamber Threshold change!
  • Tuesday

    
    

    Monday

  • 242988 09/08/08 23:00 Jin Huang Replay Dir Built for Transversity
  • 242987 09/08/08 20:41 adev BigBite WireChamber Threshold change!
  • 242982 09/08/08 20:20 adev BigBite WireChamber Threshold change!
  • 242976 09/08/08 20:10 adev BigBite WireChamber Threshold change!
  • 242974 09/08/08 19:59 adev BigBite WireChamber Measured Voltage change!
  • 242944 09/08/08 13:25 adev BigBite WireChamber HVGUI Start!
  • 242943 09/08/08 11:13 V. Sulkosky Replaced 1877 TDC in ROC 3, Slot 16
  • 242940 09/08/08 09:29 Jack Check valve on right argon cylinder cracked
  • Sunday

  • 242886 09/07/08 00:06 adev BigBite WireChamber Measured Voltage change!
  • Saturday

  • 242883 09/06/08 23:56 adev BigBite WireChamber Measured Voltage change!
  • Friday

  • 242877 09/05/08 14:47 Jin Huang Target: Spin Signal Recovered
  • 242876 09/05/08 14:23 Jin Huang Target: Spin Signal Temporarily Out
  • 242875 09/05/08 14:00 adev BigBite WireChamber Measured Voltage change!
  • 242874 09/05/08 13:50 adev BigBite WireChamber Measured Voltage change!
  • 242871 09/05/08 11:09 R. Michaels L-HRS DAQ problems fixed, some issues remain
  • Thursday

  • 242866 09/04/08 17:22 Kalyan Scaler gates ready
  • Wednesday

  • 242842 09/03/08 19:12 adev BigBite WireChamber HVGUI Start!
  • 242836 09/03/08 17:10 Chiranjib HV back to the original values
  • 242834 09/03/08 17:07 Chiranjib NO ADC histograms
  • 242829 09/03/08 16:44 Chiranjib HV change
  • 242826 09/03/08 16:26 Jin Huang Target: Simulated Spin Signal to 30s/flip
  • 242822 09/03/08 12:08 R. Michaels L-HRS scalers ready (but waiting for target state)
  • Tuesday

  • 242818 09/02/08 23:47 Jin Huang Target: Spin Signal is Back
  • 242797 09/02/08 17:17 Moffit BBDAQ: ROC5 changes. Removed software limit on pulser prescale.
  • 242792 09/02/08 15:31 fansler Q2/Q3 LN2 readbacks
  • Monday